A constant force on 40n act on a body initially at rest give it an acceleration of 0.1mls for 4seconds.calculate the work done by the force


1
Expert's answer
2022-03-04T13:23:04-0500

During this time the body will cover the following distance:


"d = \dfrac{at^2}{2}"

where "a" is the acceleration, and "t" is time.

By definition, the work is the product of the force and distance:


"W = Fd = \dfrac{Fat^2}{2}= \dfrac{40\cdot0.1\cdot 4^2 }{2} = 32J"

Answer. 32J.